Aniyah2:

Ally surveyed 5 classrooms. She wanted to know the number of students in each class who love painting. The data she collected from the 5 classrooms is below. 10, 7, 9, 11, 8 The value that could be the best measure of the center of this data is

Aniyah2:

TutorJohnfoster:

The value that could be the best measure of the center of the data is the mean. mean=\[\frac{Total\space of all\space items }{ Number\space of items }\] \[\frac{ 10+7+9+11+8 }{ 5 }\] \[\frac{ 45 }{ 5 }\] 9